Four years ago, Tom's age was 3/4 of the age he will be in 6 years. How old is he now?
The equation I tried was x - 4 = (3/4) (x + 6). I know the parentheses means multiplication, I just have never seen fractions inside them.

Answer by TimothyLamb(4379)

(t - 4) = 3/4(t + 6)
t - 4 = (3/4)t + 18/4
t - (3/4)t = 18/4 + 16/4
(4/4)t - (3/4)t = 34/4
(1/4)t = 34/4
Answer:
t = 34
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php